OpenStudy (anonymous):

What is the remainder of (x3 – 6x – 9x + 3) ÷ (x - 3) A. -51 B. -51 x - 3 C. -17 x D. -17 x - 3

OpenStudy (raden):

use the remainder theorem : if f(x) divided by (x-3), then f(3) as remainder so, what's (x^3 – 6x^2 – 9x + 3) for x=3 ? (sorry, i assumed u have typo it should -6x^2, right ?) it is 3^3 - 6(3)^2 - 9(3) + 3 = .... ?

OpenStudy (anonymous):

what i have there is the right one

OpenStudy (raden):

try this mommy :) 3^3 - 6(3)^2 - 9(3) + 3 = ?

OpenStudy (anonymous):

Thank for trying to help but I really do not get any of this

OpenStudy (anonymous):

I'm doing this all online and it is very hard to try and break it down and learn how to do all this

OpenStudy (raden):

well, i will figure out of it 3^3 - 6(3)^2 - 9(3) + 3 = 27 - 6(9) - 27 + 3 = 27 - 54 - 27 + 3 = -54 + 3 = - 51

OpenStudy (raden):

it is A, mommy :)